OpenStudy (anonymous):

simplify expression 3sqrt 64(x+5)^3

OpenStudy (solomonzelman):

\[3 \sqrt{64(x+5)^3}\] this?

OpenStudy (solomonzelman):

I'll assume that: \[3 \times \sqrt{8^2 \times (x+5)^2 \times (x+5)}\]\[3 \times 8 \times (x+5) \sqrt{(x+5)}\]\[(24x+120)\sqrt{x+5}\]\[24x \sqrt{x+5}+120\sqrt{x+5}\]